# Timo: Towards Better Temporal Reasoning for Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success in various reasoning tasks (Zhao et al., 2023; Chang et al., 2023), such as mathematical, commonsense, and symbolic reasoning. Despite these advances, LLMs face significant challenges in temporal reasoning (Chen et al., 2021; Tan et al., 2023a), which is crucial in human perception. Compared to other reasoning tasks that focus solely on one specific reasoning ability, temporal reasoning is an integrated task that requires arithmetic (Zhu et al., 2023a), logic (Mishra et al., 2022a) and world knowledge (Wei et al., 2022).
@@ 3 Self-critic Temporal Task Optimization
where $\text{align}(y_{i}^{n},g_{i})$ is a function that returns true if the response $y_{i}^{n}$ aligns with the golden label $g_{i}$, and false otherwise. Inspired by the LLM-as-a-Judge mechanism (Zheng et al., 2023; Yuan et al., 2024), we utilize mathematical model $L$ directly as a reward model to identify high-quality response pairs. Notably, we introduce a novel hierarchical scoring method, which is specifically designed for evaluating responses to temporal tasks and contains five key aspects: (1) relevance and basic temporal reasoning; (2) understanding of temporal aspects; (3) application of internal temporal knowledge; (4) direct and well-organized addressing of the question; (5) insightfulness and advanced reasoning. To choose the higher quality pair from the correct set $\mathbf{R}_{i}^{+}$ and wrong set $\mathbf{R}_{i}^{-}$, we prioritize the response that utilizes the model’s temporal reasoning to the fullest extent. The criteria for our evaluation prompts are illustrated in Figure 10 and  11. For each criterion a response meets, a point is awarded. We prompt the model $L$ to assign a score $r_{i}^{n}\in[0,5]$ to each response $y_{i}^{n}$, quantifying its quality across the above dimensions.
@@ 4.1 Experimental Setup
We conduct a comprehensive evaluation across all temporal reasoning tasks, encompassing a total of 38 tasks. Following  Tan et al. (2023a), we assess the model performance on 100 examples for each task, amounting to a total of 3,800 instances. Consistent with prior work (Hendrycks et al., 2020), we evaluate the model’s temporal abilities under the 5-shot setting and utilize greedy decoding (i.e., temperature = 0) for generating model’s responses. We extract the prediction from the response and calculate the accuracy of each subtask.
@@ 7 Related Work
Preference optimization approaches typically involve training a fixed reward model based on preference data, and then utilizing the reward model to train via reinforcement learning (RL) (Schulman et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Bai et al., 2022). To simplify this process, Direct Preference Optimization (DPO) (Rafailov et al., 2023) is introduced to avoid training the reward model entirely, and instead directly train the LLM using preference pairs. Building on this approach, recent works explore automatic optimization and self-correction in LLMs (Pan et al., 2023; Ji et al., 2024). This involves two key steps: instructing LLMs to self-generate their training dataset (Wang et al., 2023; Taori et al., 2023; Tunstall et al., 2023) and serving LLMs as reward models (Fernandes et al., 2023; Saha et al., 2023; Dubois et al., 2024) to select high-quality data. The self-generated data optimization enables models to iteratively improve their performance through a self-rewarding mechanism (Yuan et al., 2024). Inspired by the above works, we introduce a self-critic temporal optimization method that leverages the inherent capabilities of the model itself to achieve significant improvements in all temporal tasks.
@@ 8 Conclusion
Appendix A Rewarding Prompt
Our rewarding prompts are shown in Figure 10 and  11.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success in various reasoning tasks (Zhou et al., 2022; Zhao et al., 2023; Chang et al., 2023), such as mathematical, commonsense, and symbolic reasoning. Despite these advances, LLMs face significant challenges in temporal reasoning (Chen et al., 2021; Tan et al., 2023a), which is crucial in human perception. Compared to other reasoning tasks that focus solely on one specific reasoning ability, temporal reasoning is an integrated task that requires arithmetic (Zhu et al., 2023a), logic (Mishra et al., 2022a) and world knowledge (Wei et al., 2022).
@@ 3 Self-critic Temporal Task Optimization
where $\text{align}(y_{i}^{n},g_{i})$ is a function that returns true if the response $y_{i}^{n}$ aligns with the golden label $g_{i}$, and false otherwise. Inspired by the LLM-as-a-Judge mechanism (Zheng et al., 2023; Yuan et al., 2024; Qu et al., 2024b), we utilize mathematical model $L$ directly as a reward model to identify high-quality response pairs. Notably, we introduce a novel hierarchical scoring method, which is specifically designed for evaluating responses to temporal tasks and contains five key aspects: (1) relevance and basic temporal reasoning; (2) understanding of temporal aspects; (3) application of internal temporal knowledge; (4) direct and well-organized addressing of the question; (5) insightfulness and advanced reasoning. To choose the higher quality pair from the correct set $\mathbf{R}_{i}^{+}$ and wrong set $\mathbf{R}_{i}^{-}$, we prioritize the response that utilizes the model’s temporal reasoning to the fullest extent. The criteria for our evaluation prompts are illustrated in Figure 10 and  11. For each criterion a response meets, a point is awarded. We prompt the model $L$ to assign a score $r_{i}^{n}\in[0,5]$ to each response $y_{i}^{n}$, quantifying its quality across the above dimensions.
@@ 4.1 Experimental Setup
We conduct a comprehensive evaluation across all temporal reasoning tasks, encompassing a total of 38 tasks. Following  Tan et al. (2023a), we assess the model performance on 100 examples for each task, amounting to a total of 3,800 instances. Consistent with prior work (Qu et al., 2024a; Xia et al., 2024), we evaluate the model’s temporal abilities under the 5-shot setting and utilize greedy decoding (i.e., temperature = 0) for generating model’s responses. We extract the prediction from the response and calculate the accuracy of each subtask.
@@ 7 Related Work
Preference optimization approaches typically involve training a fixed reward model based on preference data, and then utilizing the reward model to train via reinforcement learning (RL) (Schulman et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Bai et al., 2022). To simplify this process, Direct Preference Optimization (DPO) (Rafailov et al., 2023) is introduced to avoid training the reward model entirely, and instead directly train the LLM using preference pairs. Building on this approach, recent works explore automatic optimization and self-correction in LLMs (Pan et al., 2023; Ji et al., 2024). This involves two key steps: instructing LLMs to self-generate their training dataset (Wang et al., 2023; Taori et al., 2023; Tunstall et al., 2023; Liu et al., ) and serving LLMs as reward models (Fernandes et al., 2023; Saha et al., 2023; Dubois et al., 2024) to select high-quality data. The self-generated data optimization enables models to iteratively improve their performance through a self-rewarding mechanism (Yuan et al., 2024). Inspired by the above works, we introduce a self-critic temporal optimization method that leverages the inherent capabilities of the model itself to achieve significant improvements in all temporal tasks.
@@ 8 Conclusion
Acknowledgement
We want to thank all the anonymous reviewers for their valuable comments. We are also grateful to Wei Liu for his insightful suggestions during the mathematical reasoning experiments. This work was supported by the National Science Foundation of China (NSFC No. 62206194), the Priority Academic Program Development of Jiangsu Higher Education Institutions, the Natural Science Foundation of Jiangsu Province, China (Grant No. BK20220488), and Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
Appendix A Prompt
Our rewarding prompts are shown in Figure 10 and  11. The prompts for different temporal tasks can be found in our public GitHub repository: https://github.com/zhaochen0110/Timo.
@@ Appendix C Iterative Optimization Study
Appendix D Validating Timo on LLaMA3-8B
To further validate the effectiveness of Timo in enhancing temporal reasoning across different LLMs, we conducted additional experiments using the LLaMA3-8B model. The results are shown in Table 6. Compared to vanilla LLaMA3-8B, Timo shows an average improvement of 5.1 scores, with 1.2 scores in math-related tasks and 9 scores in time-related tasks. These consistent improvements across both the LLaMA2 and LLaMA3 series demonstrate Timo’s strong generalization capabilities across different model series, enhancing its applicability and effectiveness in diverse settings.
Math-time Pure-time Average
LLaMA3-8B 81.4 79.6 80.5
+Timo 82.6 88.6 85.6
Table 6: Performance Comparison of LLaMA3-8B with and without Timo enhancement
Appendix E Evaluating the Impact of Math LLM on Temporal Reasoning
Existing work on weak-to-strong generalization suggests that distilling data from a weaker or equivalent LLM benefits a stronger LLM (Burns et al., 2023). To address concerns regarding the influence of the LLM-as-Judge framework compared to the use of a specialized math LLM, we conducted experiments using vanilla LLaMA2-7B and LLaMA2-7B-chat, representing general SFT LLaMA models. As presented in Table 7, our results demonstrate that incorporating a math LLM yields significant improvements in temporal reasoning tasks. Specifically, the math LLM outperforms the vanilla LLaMA2-7B and LLaMA2-7B-chat models by an average of 3.6 and 7 scores, respectively. The performance gains are especially notable in math-related tasks, where the math LLM achieves scores 5.5 and 10.5 scores higher than those of the other two models. These results indicate that the math LLM component is crucial for enhancing temporal reasoning capabilities, outperforming the self-critic temporal optimization (i.e., LLM-as-Judge) framework alone. The results indicate that math-specific training plays a pivotal role in reasoning over time, confirming the value of specialized LLMs in complex reasoning tasks.
Math-time Pure-time Average
Timo (LLaMA2-7B) 58.4 79.7 69.1
Timo (LLaMA2-7B-chat) 53.4 78.1 65.7
Timo (MathLLaMA-7B) 63.9 81.5 72.7
Table 7: Comparison of temporal reasoning performance across different based LLM, with Timo applied for temporal optimization on LLaMA2-7B, LLaMA2-7B-chat, and MathLLaMA-7B.
Appendix F Further Evaluation of Timo on Temporal Reasoning Datasets
To further assess Timo’s improvements in temporal reasoning, we extended our evaluation to additional temporal reasoning datasets, i.e., MCTACO (Zhou et al., 2019) and TempReason (Tan et al., 2023a). These datasets were selected to validate Timo’s effectiveness across a broader range of temporal reasoning tasks.
  • MCTACO: This dataset evaluates a wide range of commonsense knowledge related to events, including the duration, frequency, order, stationary nature, and typical timing of events.
  • TempReason: This dataset emphasizes implicit temporal reasoning in structured facts, focusing on both event-time reasoning and event-event reasoning.
The results are shown in Table 8. Timo achieves scores 6.2 and 15.5 points higher than LLaMA2-7B and WizardMath-7B on the TempReason task. Additionally, Timo surpasses MAmmoTH-7B by 19.3 points on the MCTACO task. These results indicate that Timo excels across various temporal reasoning datasets, demonstrating its robust general temporal reasoning abilities. In future work, We will further explore the generalization of Timo across more reasoning tasks, such as commonsense reasoning (Sakaguchi et al., 2021), and composition relations reasoning (Zhao & Zhang, 2024).
MCTACO TempReason
LLaMA2-7B 50.3 46.6
MAmmoTH-7B 37.0 52.8
WizardMath-7B 12.7 37.3
Timo-7B 56.3 52.8
Table 8: Results on the MCTACO and TempReason datasets. Timo-7B outperforms its counterparts, demonstrating superior general temporal reasoning abilities.